Currently, the firm has excess cash of $1,500, other assets of $5,000, and equity valued at $4,000. The firm has 300 shares of stock outstanding and net income of $380. What will the new earnings per share (EPS) be if the firm uses 20 percent of its excess cash to complete a stock repurchase? Please assume that the market value of the firm equal to its book value.

Group of answer choices $1.27 $1.36 $1.68 $1.83 None of above is correct.
Answer #1

Value per share = Value of Equity/Number of shares

= 4000/300

= $13.33 per share

Number of shares bought back = 1500*20%/13.33 = 22.51 shares

EPS after repurchase = Net Income/Number of shares

= 380/(300-22.51)

= $1.3694 per share

i.e. $1.36 per share approx.